A bicycler traveling 12 mph passes a runner traveling 9 mph. If the runner started at noon, and the bicycler at 12:30, how far from the start point did they meet?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

runner 9 mph
bicycler 12 mph
runner 12:00
bicycler 12:30
Difference in time= 00:30 0.5
Runner will have covered 4.5 miles before bicycler starts
catch up distance= 4.5 miles
catch up speed = 12 -9 mph
catch up speed = 3 mph
Catchup time = catchup distance/catch up speed
catch up time= 4.5 / 3
catch up time= 1.5 hours
bicycler speed = 12 mph
time = 1.5
12*1.5= 18 miles

Answer by ikleyn(53938)

.
A bicycler traveling 12 mph passes a runner traveling 9 mph.
If the runner started at noon, and the bicycler at 12:30, how far from the start point did they meet?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

When the bicycler started at 12:30, the runner was 9*0.5 = 4.5 miles ahead.


The approaching speed was  12-9 = 3 mph,  so the bicycler catch up the runner in  4.5/3 = 1.5 hours.


During this time, the bicyclist covered 1.5*12 = 18 miles from the start.


So, the distance from the start to the meeting point was 18 km.    ANSWER

Solved.